# Partition functions of determinantal point processes on polarized Kähler manifolds

### v3

@@ Appendix A Explicit computations on the generalized Liouville action
Remark 6.
Proof 1 shows that for any smooth path $\varphi_{t}$ in $\mathcal{K}_{\omega}$ joining 00 and $\varphi$, we have
$\int_{M}-iBC(\operatorname{Td}_{2};\omega_{\varphi},\omega)\frac{\omega^{n-1}}{(n-1)!}+\operatorname{Td}_{2}(R_{\varphi})\frac{-1}{(n-1)!}\sum^{n-2}_{s=0}\varphi\omega_{\varphi}^{s}\wedge\omega^{n-2-s}$ (A.40)
[…]
without polarization assumption on $[\omega]$.
From Proof 2, we can read off the second variation formula of $S_{2}$.
Proposition A.1.
$\;-\frac{1}{2}\Delta\dot{\varphi}\langle i\partial\bar{\partial}\dot{\varphi},\operatorname{Ric}\rangle\frac{\omega^{n}}{n!}+\frac{1}{12}\left(g^{i\bar{q}}g^{p\bar{j}}g^{r\bar{l}}g^{k\bar{s}}R_{p\bar{q}r\bar{s}}\partial_{i}\partial_{\bar{j}}\dot{\varphi}\partial_{k}\partial_{\bar{l}}\dot{\varphi}\right)\frac{\omega^{n}}{n!}.$ (A.42)

### v5

@@ Appendix A Explicit computations for the third coefficient
Proof 1 shows that for any smooth path $\varphi_{t}$ in $\mathcal{K}_{\omega}$ joining 0 and $\varphi$, we have
$\int_{M}-iBC(\operatorname{Td}_{2};\omega_{\varphi},\omega)\frac{\omega^{n-1}}{(n-1)!}+\operatorname{Td}_{2}(R_{\varphi})\frac{-1}{(n-1)!}\sum^{n-2}_{s=0}\varphi\omega_{\varphi}^{s}\wedge\omega^{n-2-s}$
$=\int_{0}^{1}dt\int_{M}\dot{\varphi_{t}}\left(\frac{1}{6}\Delta_{t}S_{t}-\frac{1}{24}\left(|R_{t}|^{2}-4|\operatorname{Ric}_{t}|^{2}+3S_{t}^{2}\right)\right)\frac{\omega_{t}^{n}}{n!},$
without polarization assumption on $\omega]$.
From Proof 2, we can read off the second variation formula of $S_{2}$, Proposition 1.